OpenStudy (anonymous):

Prove that the sum of i*2^(i-1) from i=1 to n equals (n-1)*2^n using mathematical induction, where n is all natural numbers.

OpenStudy (anonymous):

Am I doing something stupid, because for some reason I'm getting different answers for the right and left sides for n = 1

OpenStudy (freckles):

\[\sum_{i=1}^{n}i \cdot 2^{i-1}=(n-1)2^n\]

OpenStudy (anonymous):

^ Yes, that's correct.

OpenStudy (freckles):

n=1 we have \[1 \cdot 2^{1-1}=1 \\ (1-1)2^{1}=0\] you are right two different things

OpenStudy (anonymous):

My teacher's the one who made this worksheet... maybe he made a mistake. I'll ask him tmw thanks.

OpenStudy (freckles):

i wonder if it works for some restriction on the natural numbers...

OpenStudy (freckles):

does it work for n=2?

OpenStudy (freckles):

\[1 \cdot 2^{1-1}+2 \cdot 2^{2-1}=1+4=5 \\ (2-1)2^{2}=1(4)=4\]

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

He probably just made a mistake... I hope, our test is on Wednesday, ha ha.

OpenStudy (freckles):

@ganeshie8 do you think you see the mistake

OpenStudy (freckles):

I wonder if we can correct the problem and then do the problem because you know you want to be prepared for this test in case he puts correct problems on there

OpenStudy (aum):

Seems to be off by 1. Adding a 1 to RHS might do it.

OpenStudy (anonymous):

You know, he did go on a rant the other day about how we need to know the struggle of coming up with the formula the other day... I'll give it a shot.

ganeshie8 (ganeshie8):

yeah = (n-1)*2^n + 1

OpenStudy (freckles):

@aum noticed something all the things we got differ by 1

OpenStudy (anonymous):

Right of course.

OpenStudy (freckles):

Do you want to try to do the problem now that has been corrected before we assist?

OpenStudy (anonymous):

Yup, I'll probably be able to do it now. Thanks :D

OpenStudy (xapproachesinfinity):

i also agree it should be (n-1)*2^n+1

OpenStudy (anonymous):

Yup, induction proof works now. Thanks everyone.

OpenStudy (xapproachesinfinity):

seems to me n=3 doesn't work!

OpenStudy (freckles):

\[1 \cdot 2^{1-1}+2 \cdot 2^{2-1}+3 \cdot 2^{3-1}=1+4+12=17 \\ (3-1)2^{3}=2(8)=16\]

OpenStudy (freckles):

oh and of course add 1 to the bottom there

OpenStudy (freckles):

unless i didn't plug in correctly i'm kinda feeling jittery a little so my screen is shaking in mind

OpenStudy (xapproachesinfinity):

it is working! i just miss calculated

ganeshie8 (ganeshie8):

\[\large \sum\limits_{i=0}^nx^i = \dfrac{x^{n+1}-1}{x-1}\]

ganeshie8 (ganeshie8):

for a non induction version of the proof, this trick works i think differentiate both sides with respect to x and plugin in x=2

OpenStudy (anonymous):

This is what I did (ignore the stuff in the corner and at the bottom)

ganeshie8 (ganeshie8):

im not seeing the base case n=1 ?

OpenStudy (aum):

Also the "i" on the LHS should remain as "i". Only the n should be replaced with "k".

ganeshie8 (ganeshie8):

also the index variable on left hand side is dummy, don't touch it

OpenStudy (anonymous):

Dummy?

ganeshie8 (ganeshie8):

\[\sum_{i=1}^{n}i \cdot 2^{i-1}=(n-1)2^n\] \[\sum_{j=1}^{n}j \cdot 2^{j-1}=(n-1)2^n\] \[\sum_{\spadesuit =1}^{n}\spadesuit \cdot 2^{\spadesuit -1}=(n-1)2^n\]

ganeshie8 (ganeshie8):

the sum doesn't depend on what index variable you choose

ganeshie8 (ganeshie8):

thats the reason we call it dummy

ganeshie8 (ganeshie8):

** add +1 on right hand side

OpenStudy (anonymous):

Okay got it. Thanks.

ganeshie8 (ganeshie8):

for an induction proof, the base case is very important. you can't get away with the proof without showing the statement works for n=1

ganeshie8 (ganeshie8):

Proposition\(\large P(n)\) : \(\sum\limits_{i=1}^ni\cdot2^{i-1} = (n-1)2^n+1\) Step1 (Basis of the induction) : \( P(1) = \sum\limits_{i=1}^1i\cdot2^{i-1} = 1\cdot 2^0 = (1-1)2^1 + 1 ~\color{red}{\checkmark } \) so the given statement works for n=1

ganeshie8 (ganeshie8):

Step2 (Assumption) : assume \(\large P(k)\) is true \[\large \sum\limits_{i=1}^ki\cdot 2^{i-1} = (k-1)2^k+1 \]

ganeshie8 (ganeshie8):

Step3 (Induction step) : prove \(\large P(k+1)\) is true \[\large \cdots \] \[\large \cdots \]

ganeshie8 (ganeshie8):

thats the general structure you need to show in any induction proof

ganeshie8 (ganeshie8):

all 3 steps are necessary for the proof to work

OpenStudy (anonymous):

Alright. I just cut corners because I knew it worked in my head/here and my teacher never checks the homework anyways. I'd write the whole thing on a test.

ganeshie8 (ganeshie8):

that sounds good, your work for P(k+1) looks nice :)

OpenStudy (anonymous):

Thanks, but I know it was a simple question, I'm 100% sure that it's gonna be ten times harder on my test... my teacher loves making questions so hard half the class fails, ha ha. but it preps us for uni I guess.